Grade 11Mechanics

The equation of the path of the projectile is y=0.5x-0.04x^2.The initial speed of projectile is
A)10m/s
B)15m/s
C)12.5m/s
D)None

jiten
8 Years agoGrade 11

2 Answers

Eshan
8 Years ago
Dear student,

The equation of path of a projectile is given as-

y=xtan\theta-\dfrac{1}{2}\dfrac{gx^2}{u^2cos^2\theta}
From the given equation,y=0.5x-0.04x^2

tan\theta=0.5
\implies cos\theta=\dfrac{2}{\sqrt{5}}

and\dfrac{g}{2u^2cos^2\theta}=0.04
\implies u=\sqrt{\dfrac{g}{2(0.04)cos^2\theta}}=12.5m/s
Khimraj
8 Years ago
from the equation of projectile
y = xtan\Theta  – ½(gx2)/(u2cos2\Theta)
tan\Theta = ½ 
so cos\Theta  = 2/\sqrt{5}
and
g/2u2cos2\Theta = 0.04
SO u = 12.5 m/s.
Hope it clears.